非線形 SCHRÖDINGER 方程式の修正された散乱状態の存在について

林 仲夫

次の非線形 Schrödinger 方程式の修正された散乱状態の存在について考える.

\{ \begin{matrix}iu_{t} + \frac{1}{2}u_{xx} = a (t) F (u, u_{x}) , x ∈ R, t > 0, \\ u (0, x) = εu_{0} (x) , x ∈ R, \end{matrix}

ここで係数 a (t) は次の条件を満足すると仮定し ε > 0 は十分小さいとする.

|a (t)| ≤ C (1 + |t|)^{1−δ} , δ ∈ (0, 1) .

また非線形項は

F (u, u_{x}) = λ_{1} |u|^{2} u + iλ_{2} |u|^{2} u_{x} + iλ_{3}u_{x}^{2}\overline{u} + λ_{4} |u_{x}|^{2} u
+λ_{5}\overline{u}u_{x}^{2} + iλ_{6} |u_{x}|^{2} u_{x}

とし係数は λ_{1}, λ_{6} ∈ R, λ_{2}, λ_{3}, λ_{4}, λ_{5} ∈ C, λ_{2} − λ_{3} ∈ R, λ_{4} − λ_{5} ∈ R を満足すると仮定する. 文献 [5] において初期値が指数関数的に減衰している場合に次のような û_{+} の存在と解の漸近的振る舞いを示した.

u (t, x) = MDû_{+} exp (−i ∫_{0}^{t} \frac{a (τ) dτ}{τ + 1} b \left|{ \atop û_{+}}\right|^{2} + O (1 + t^{1−2δ}))
+O (t^{−1/2−δ})

ここで

b(ξ) = λ_{1} − (λ_{2} − λ_{3}) ξ + (λ_{4} − λ_{5}) ξ^{2} − λ_{6}ξ^{3},

(D (t) ϕ) (x) = (\frac{1}{it})^{1/2} ϕ (\frac{x}{t}) , M = exp (ix^{2}/2t)

とする. δ ∈ (\frac{1}{2}, 1) の場合は次のような漸近的振る舞いが得られる.

u (t, x) = MDû_{+} exp (−i ∫_{0}^{t} \frac{a (τ) dτ}{τ + 1} b \left|{ \atop û_{+}}\right|^{2}) + O (t^{1−2δ})

∥u (t) − exp (−i ∫_{0}^{t} \frac{a (τ) dτ}{τ + 1} b (\frac{·}{t}) |û_{+} (\frac{·}{t})|^{2}) U (t) u_{+}∥_{L^{2}} ≤ Ct^{1−2δ}

この講演では δ ∈ (1/2, 1) について考える.

記号及び関数空間. Fourier 変換を Fϕ あるいは ϕ̂ で記述し

(Fϕ) (ξ) = \frac{1}{\sqrt{2π}} ∫ e^{−ixξ}ϕ (x) dx

で定義する. 通常の Lebesgue space を

L^{p} = {ϕ ∈ S′; ∥ϕ∥_{p} < ∞}

1This is a joint work with P.I.Naumkin

Date: 4 月 28 日, 2001.

で定義し重み付の Sobolev 空間を

H^{m,s} = {ϕ ∈ L^{2}; ∥ϕ∥_{m,s} = ∥ ⟨x⟩^{s} ⟨i∂_{x}⟩^{m} ϕ∥_{2} < ∞} , m, s ∈ R^{+},

とする. ただしここで ⟨x⟩ = \sqrt{1 + x^{2}}. とである. また解析関数の空間を次のように定義する.

H_{σ}^{m,s} = {ϕ ∈ L^{2}; ∥ϕ∥_{H_{σ}^{m,s}} = ∥e^{σ|p|} ⟨p⟩^{m} ⟨i∂_{p}⟩^{s} ϕ̂ (p) ∥_{2} < ∞}, m, s ∈ R^{+},

次の結果を得た.

Theorem 0.1. We assume that δ ∈ (0, 1) and the initial data are such that e^{β⟨x⟩}u_{0} ∈ H^{3,5/2} with β > 0. Then there exists a unique modified final state W ∈ H_{β/2}^{5/2,3}, and

unique real-valued phase function G (t) ∈ C ([1, ∞) ; H_{β/2}^{2,3}) , such that

∥u(t) − MDW e^{iG(t)}∥_{p} ≤ Cεt^{− \frac{1}{2}(1− \frac{2}{p})−δ} (0.1)

for all t ≥ 1, where 2 ≤ p ≤ ∞.

Remark 0.1. From the proof of Theorem 0.1 we can see that G (t) = 0 in the case b = 0, i.e. λ_{1} = 0, λ_{2} − λ_{3} = 0,

λ_{4} − λ_{5} = 0, λ_{6} = 0. Therefore this exceptional nonlinearity appears to be super critical since the usual scattering states exist. In the opposite case estimate (0.1) shows the existence of the modifies scattering states in the whole region δ ∈ (0, 1) .

証明の方法は文献 [2] 及び [3] において用いられた位相関数の高次の近似と以前 [5] で用いた方法を利用する.

References

[1] J. Ginibre and G. Velo, Long range scattering and modified wave operators for some Hartree type equations I, Rev. Math. Phys., to appear.

[2] J. Ginibre and G. Velo, Long range scattering and modified wave operators for some Hartree type equations II, Ann. Henri Poincaré, 1 (2000), pp. 753-800.

[3] J. Ginibre and G. Velo, Long range scattering and modified wave operators for some Hartree type equations III, Gevrey spaces and low dimensions, preprint (2000).

[4] N. Hayashi, P.I. Naumkin and P.N. Pipolo, Analytic smoothing effects for some derivative nonlinear Schrödinger equations, Tsukuba Math. J., 24 (2000), pp. 21-34.

[5] N. Hayashi, P.I. Naumkin and Y. Yamazaki, Large time behavior of small solutions to subcritical derivative nonlinear Schrödinger equations, Proc. Amer. Math. Soc., to appear.

大阪大学

   E-mail address: nhayashi@math.wani.osaka-u.ac.jp